# The Rise of Parameter Specialization for Knowledge Storage in Large Language Models

### v1

@@ 1 Introduction
An increasing number of powerful large language models (LLMs) have emerged in recent years (Touvron et al., 2023a; Achiam et al., 2023; Groeneveld et al., 2024; Bai et al., 2023; Team, 2025), often demonstrating remarkable capabilities across various benchmarks and tests (Hendrycks et al., 2021a; Chen et al., 2021a; Cobbe et al., 2021). Thanks to the large parameter space, they have shown an exceptional ability to encode vast amounts of knowledge within their parameters, enabling superior performance on knowledge-intensive tasks (Hendrycks et al., 2021a; Zhang et al., 2023).
To understand the internal mechanism of knowledge storage, many studies have been conducted. For example, Geva et al. (2021b) interprets the MLP layers of the transformer architecture (Vaswani et al., 2017) as key-value memories, where the factual knowledge encoded in the weights is retrieved and transmitted to the output layer during inference (Geva et al., 2023; Meng et al., 2022; Yu et al., 2024). Furthermore, researchers have observed that, in the final layer of the MLP, each vector in that value matrix can act as a fundamental unit of knowledge storage (Geva et al., 2022a, b). However, there has been limited research on how to better store and compress knowledge within constrained model parameters to enable more effective utilization of that knowledge by the model.
In this work, we investigate the relationship between language models’ knowledge storage patterns and their performance. To identify parameters associated with specific knowledge concepts, we analyze consistently activated parameters in MLP layers when the model processes questions related to the particular knowledge concept. Building on the key-value interpretation of the MLP by Geva et al. (2021b), which treats the up-projection matrix as the key and the down-projection matrix as the value (i.e., stored knowledge), we extract the intermediate representations between these two matrices and treat their absolute value as the activation of corresponding parameters. To support empirical analysis, we construct a new encyclopedic knowledge benchmark based on Wikipedia, covering knowledge concepts with varying frequencies. We then apply the knowledge parameter identification method to 20 open-source LLMs across a wide range of model families, enabling us to explore correlations between knowledge storage patterns and overall model performance.
@@
Motivated by this observation, we further conduct four sets of controlled experiments, each involving continued training on the Llama2-7B (Touvron et al., 2023a) and Qwen2-7B (Yang et al., 2024) models with new knowledge respectively, to validate the strong causal relationship between improved parameter specialization and enhanced performance of the models on knowledge tasks. Overall, the experiments reveal that encoding similar knowledge into the same parameter vectors better aligns with the model’s internal knowledge retrieval mechanism. This approach helps the model utilize knowledge more efficiently, improves knowledge compression, and reduces hallucination generation.
@@ 2 Related Work
Studying how knowledge is stored and utilized in LLMs has been an important area in the research of LLM interpretability (Meng et al., 2022; Geva et al., 2021b; Sukhbaatar et al., 2015; Geva et al., 2023). Recent studies have shown that MLPs are the primary and crucial components for storing factual knowledge and associations in transformer-based language models (Geva et al., 2022b; Dar et al., 2023). They can be conceptualized as key-value memories (Geva et al., 2021b), where the factual knowledge encoded in the MLP weights is recalled and transmitted to the output layer during inference (Geva et al., 2023; Meng et al., 2022; Yu et al., 2024). Additionally, researchers have found that in the final layer of the MLP, each vector in the value matrix can serve as a fundamental unit for storing knowledge (Geva et al., 2022a, b). They have also verified that by directly manipulating or disrupting these parameter vectors, specific knowledge can be edited or unlearned (Hong et al., 2024a, b; Meng et al., 2022), leading to changes in the model’s responses.
@@
Elhage et al. (2022); Olah (2023) propose the concept of Knowledge Superposition. It refers to an inevitable phenomenon in neural network models, especially large language models, during training and data memorization: since the number of data features greatly exceeds the number of parameters in the model, each parameter does not have a simple one-to-one mapping with the data features or knowledge. Neurons are often involved with multiple data features simultaneously. In our work, we treat each vector in the last layer of MLP as a basic unit for storing knowledge and investigate the superposition of knowledge within these vectors.
@@ 3.1 Preliminary
In transformer-based language models, the MLP is a crucial component for storing the model’s factual knowledge, and its sub-layers can be viewed as key-value memories (Geva et al., 2021b). To be specific, the first layer***In most decoder-only models, such as GPT-2 (Radford et al., 2019) and GPT-J (Chen et al., 2021b), the MLP component consists of two layers, whereas in LLaMA (Touvron et al., 2023b), it comprises three layers. However, we can still regard LLaMA’s first two layers collectively as the key matrices, with their output representing the coefficient scores. of MLP sublayers can be viewed as a matrix $W_{K}$ formed by key vectors $\{\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{n}\}$, used to capture a set of patterns in the input sequence, and ultimately outputting the coefficient scores. The second layer can be viewed as a matrix $W_{V}$ formed by value vectors $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}\}$, with each value vector containing the corresponding factual knowledge.
@@ 3.2 Knowledge Vectors Masking Procedure
Therefore, given a concept, when we aim to identify which specific value vectors in the model’s MLPs are most closely related to the knowledge contained in that concept—while avoiding the masking of vectors associated with the model’s general capabilitiesThe term ”general ability” refers to the model’s fundamental skills, such as processing text inputs correctly and generating coherent outputs, rather than encoding knowledge specific to a particular concept. (Meng et al., 2022; Geva et al., 2023), i.e., determining the appropriate subset $S^{\ell}$ at each layer of the model for this concept—we will run $t$ concept-related questions and $t^{*}$ irrelevant questions on the selected model. Then we will compute the corresponding coefficients $\mathbf{m}^{\ell}$ and $\mathbf{m}^{*\ell}$, which are the averages of the coefficients for the concept-related questions and irrelevant questions, respectively, at each layer of the model. For details on the generation of concept-related and irrelevant questions, as well as the selection of $t$ and $t^{*}$, please refer to §3.4. After obtaining $\mathbf{m}^{\ell}$ and $\mathbf{m}^{*\ell}$ at each layer, we perform the computation using the following formula:
[…]
@@ 3.3 The Definition of Parameter Specialization
$\text{PSS}\stackrel{{\scriptstyle\triangle}}{{=}}\frac{|\text{General Score after surgery}-\text{Concept Specific Score after surgery}|}{\text{General Score before surgery}},$ (6)
@@ 3.4 Dataset Construction
Given that the frequency of knowledge in training datasets significantly influences a model’s ability to retain and comprehend it (Allen-Zhu & Li, 2023; Meng et al., 2022; Mallen et al., 2023), we utilize Wikipedia page views as a proxy for knowledge frequency in the models’ pre-training datasetsTo better support this point, we include experiments in §B.4 of Appendix that validate the strong correlation between concept popularity and their frequency in the pretraining data.. To this end, we calculated the page views for each concept on Wikipedia between January 1, 2010, and December 31, 2019The earliest release date of the evaluated models, such as GPT-J, is 2019. Therefore, their pretraining datasets could not include knowledge or concepts that emerged after this period. To ensure a fair evaluation across all models, any knowledge introduced post-2019, including the COVID-19 pandemic, was excluded from the benchmark. . Based on these statistics, concepts are categorized by page view frequency into three equal tiers: low-frequency (bottom 33% of the distribution), medium-frequency (middle 33%), and high-frequency (top 33%). A more detailed distribution of the categories and the corresponding example data of SpecWiki dataset are provided in Table 4 and Table 5, respectively, in the Appendix. This approximation helps estimate the likelihood of a concept’s presence in the models’ pre-training datasets and allows us to explore how knowledge at different frequency levels is stored in models and provides a more comprehensive evaluation.
@@
  • Multi-Choice Questions. Drawing inspiration from the widely used Massive Multitask Language Understanding (MMLU) dataset (Hendrycks et al., 2021b), which evaluates general knowledge across models, we similarly designed ten multiple-choice questions for each concept, ensuring the knowledge and answers could be directly found in the relevant Wikipedia articles. Specifically, we provided GPT-4o (OpenAI et al., 2024) with the appropriate Wikipedia article for each concept and instructed it to extract ten questions without overlap, along with the correct answers derived from the article’s text. Next, it was instructed to generate three additional incorrect answers, aside from the golden answer, ensuring that none of them overlapped with the correct answer to avoid confusion. The detailed prompt is available in §B.1. We also include sample multiple-choice questions and results of manual verification of the generated data in Appendix §B.2.
@@ 4.1 Experimental Setup
To provide a more comprehensive evaluation of how the degree of parameter specialization evolves across large language models, we assessed 20 open-source models from various families and sizes in the community. Specifically, we evaluated LLaMA series (Touvron et al., 2023a, b; Grattafiori et al., 2024), Qwen series (Bai et al., 2023; Yang et al., 2024; Team, 2025), Gemma series (Team et al., 2024a, b), OLMo series (Groeneveld et al., 2024; OLMo et al., 2025), Yi series (AI et al., 2025), Mistral series (Jiang et al., 2023), GPT-j-6b (Wang & Komatsuzaki, 2021), Pythia-6.9b (Biderman et al., 2023), Falcon-7b (Almazrouei et al., 2023) and Mpt-7b (Databricks, 2023). Refer to Appendix §C.1 for the implementation details of these models.
@@
Regarding the selection of model layers for masking, since the initial layers of a model typically handle fundamental capabilities like basic text processing (Meng et al., 2022; Geva et al., 2023), masking these layers could severely impair the model’s basic text generation abilities. Therefore, for all models in our study, we preserve the first 5 layers without masking and only apply vector masking operations to all subsequent layers.
@@ 4.2 Main Results
The main results for the Multiple-Choice Questions setting can be seen in Figure 2. We observe a strong correlation between the degree of Parameter Specialization (measured by PSS) and model performance on SpecWiki across 20 models, with Pearson and Spearman coefficients of 0.92 and 0.93, respectively. Models achieving better performance on SpecWiki exhibit higher Parameter Specialization Scores. Furthermore, models with higher PSS are often those released more recently (warmer color) and exhibit stronger general abilities, as measured by their MMLU performance (larger circle). The corresponding results for the Open-ended Generation setting can be found in Figure 6 in §C.2, which exhibit similar patterns and trends.
@@
Model $\textbf{Accuracy}_{\textbf{MCQ}}$ $\bm{\uparrow}$ PSS $\bm{\uparrow}$
Qwen1.5-0.5B 0.61 (±$(\pm 0.2)$ 0.019 (±$(\pm 0.01)$
Qwen1.5-1.8B 0.61 (±$(\pm 0.3)$ 0.044 (±$(\pm 0.02)$
Qwen1.5-4B 0.73 (±$(\pm 0.2)$ 0.075 (±$(\pm 0.02)$
Qwen1.5-7B 0.75 (±$(\pm 0.2)$ 0.121 (±$(\pm 0.04)$
Qwen1.5-14B 0.82 (±$(\pm 0.2)$ 0.184 (±$(\pm 0.03)$
Gemma2-2B 0.72 (±$(\pm 0.3)$ 0.057 (±$(\pm 0.02)$
Gemma2-9B 0.86 (±$(\pm 0.1)$ 0.138 (±$(\pm 0.03)$
@@ 4.4 Evolution of Parameter Specialization During Pretraining
To better investigate the development of Parameter Specialization in the model from the perspective of model training dynamics, we analyzed 10 checkpoints from the OLMo-2-1124-7B (OLMo et al., 2025) pretraining process by using our SpecWiki. The results are shown in Figure 4 below.
@@ 5 Validation of Parameter Specialization Benefits for Knowledge Tasks
Model $\textbf{Accuracy}_{\textbf{MCQ}}$ $\bm{\uparrow}$ $\textbf{Accuracy}_{\textbf{OEG}}$ $\bm{\uparrow}$ PSS $\bm{\uparrow}$ Semantic Entropy $\bm{\downarrow}$ Local Intrinsic Dimension $\bm{\downarrow}$
LLaMA2-7B 0.60 (±$(\pm 0.2)$ 0.51 (±$(\pm 0.1)$ 0.67 (±$(\pm 0.1)$ 0.67 (±$(\pm 0.1)$ 11.23 (±$(\pm 2.1)$
LLaMA2-7BFT-FV 0.63 (±$(\pm 0.3)$ 0.54 (±$(\pm 0.2)$ 0.65 (±$(\pm 0.2)$ 0.62 (±$(\pm 0.1)$ 11.12 (±$(\pm 1.4)$
LLaMA2-7BFT-PV 0.67 (±$(\pm 0.3)$ 0.59 (±$(\pm 0.2)$ 0.72 (±$(\pm 0.1)$ 0.50 (±$(\pm 0.2)$ 7.89 (±$(\pm 2.1)$
LLaMA2-7BFT-CV 0.62 (±$(\pm 0.1)$ 0.51 (±$(\pm 0.1)$ 0.63 (±$(\pm 0.2)$ 0.62 (±$(\pm 0.1)$ 11.12 (±$(\pm 1.4)$
LLaMA2-7BFT-RV 0.58 (±$(\pm 0.2)$ 0.49 (±$(\pm 0.2)$ 0.65 (±$(\pm 0.2)$ 0.65 (±$(\pm 0.2)$ 11.07(±$(\pm 2.7)$
Qwen2-7B 0.72 (±$(\pm 0.3)$ 0.63 (±$(\pm 0.1)$ 0.124 (±$(\pm 0.03)$ 0.56 (±$(\pm 0.1)$ 9.78 (±$(\pm 1.9)$
Qwen2-7BFT-FV 0.73 (±$(\pm 0.2)$ 0.67 (±$(\pm 0.2)$ 0.110 (±$(\pm 0.02)$ 0.59 (±$(\pm 0.2)$ 8.53 (±$(\pm 1.1)$
Qwen2-7BFT-PV 0.77 (±$(\pm 0.1)$ 0.70 (±$(\pm 0.1)$ 0.133 (±$(\pm 0.03)$ 0.39 (±$(\pm 0.1)$ 6.92 (±$(\pm 1.3)$
Qwen2-7BFT-CV 0.73 (±$(\pm 0.2)$ 0.65 (±$(\pm 0.2)$ 0.114 (±$(\pm 0.03)$ 0.55 (±$(\pm 0.2)$ 8.78 (±$(\pm 1.6)$
Qwen2-7BFT-RV 0.71 (±$(\pm 0.2)$ 0.63 (±$(\pm 0.1)$ 0.122 (±$(\pm 0.02)$ 0.59 (±$(\pm 0.2)$ 9.65 (±$(\pm 1.4)$
@@
In this section, we conducted four sets of controlled training experiments, each involving continued fine-tuning on the Llama2-7B-base (Touvron et al., 2023a) and Qwen2-7B-base (Yang et al., 2024) models with additional knowledge data. These experiments aim to validate the causal relationship between improved parameter specialization and enhanced model performance on knowledge tasks.

### v2

@@ 1 Introduction
An increasing number of powerful large language models (LLMs) have emerged in recent years (touvron2023llamaopenefficientfoundation; achiam2023gpt; groeneveld2024olmo; bai2023qwentechnicalreport; qwen3blog), often demonstrating remarkable capabilities across various benchmarks and tests (hendrycks2021measuring; chen2021evaluatinglargelanguagemodels; cobbe2021trainingverifierssolvemath). Thanks to the large parameter space, they have shown an exceptional ability to encode vast amounts of knowledge within their parameters, enabling superior performance on knowledge-intensive tasks (hendrycks2021measuring; zhang2023m3exam).
To understand the internal mechanism of knowledge storage, many studies have been conducted. For example, geva2021transformer interprets the MLP layers of the transformer architecture (vaswani2017attention) as key-value memories, where the factual knowledge encoded in the weights is retrieved and transmitted to the output layer during inference (geva2023dissecting; meng2022locating; yu2024mechanisms). Furthermore, researchers have observed that, in the final layer of the MLP, each vector in that value matrix can act as a fundamental unit of knowledge storage (geva-etal-2022-lm; geva-etal-2022-transformer). However, there has been limited research on how to better store and compress knowledge within constrained model parameters to enable more effective utilization of that knowledge by the model.
In this work, we investigate the relationship between language models’ knowledge storage patterns and their performance. To identify parameters associated with specific knowledge concepts, we analyze consistently activated parameters in MLP layers when the model processes questions related to the particular knowledge concept. Building on the key-value interpretation of the MLP by geva2021transformer, which treats the up-projection matrix as the key and the down-projection matrix as the value (i.e., stored knowledge), we extract the intermediate representations between these two matrices and treat their absolute value as the activation of corresponding parameters. To support empirical analysis, we construct a new encyclopedic knowledge benchmark based on Wikipedia, covering knowledge concepts with varying frequencies. We then apply the knowledge parameter identification method to 20 open-source LLMs across a wide range of model families, enabling us to explore correlations between knowledge storage patterns and overall model performance.
@@
Motivated by this observation, we further conduct four sets of controlled experiments, each involving continued training on the Llama2-7B (touvron2023llamaopenefficientfoundation) and Qwen2-7B (yang2024qwen2technicalreport) models with new knowledge respectively, to validate the strong causal relationship between improved parameter specialization and enhanced performance of the models on knowledge tasks. Overall, the experiments reveal that encoding similar knowledge into the same parameter vectors better aligns with the model’s internal knowledge retrieval mechanism. This approach helps the model utilize knowledge more efficiently, improves knowledge compression, and reduces hallucination generation.
@@ 2 Related Work
Studying how knowledge is stored and utilized in LLMs has been an important area in the research of LLM interpretability (meng2022locating; geva2021transformer; sukhbaatar2015end; geva2023dissecting). Recent studies have shown that MLPs are the primary and crucial components for storing factual knowledge and associations in transformer-based language models (geva-etal-2022-transformer; dar-etal-2023-analyzing). They can be conceptualized as key-value memories (geva2021transformer), where the factual knowledge encoded in the MLP weights is recalled and transmitted to the output layer during inference (geva2023dissecting; meng2022locating; yu2024mechanisms). Additionally, researchers have found that in the final layer of the MLP, each vector in the value matrix can serve as a fundamental unit for storing knowledge (geva-etal-2022-lm; geva-etal-2022-transformer). They have also verified that by directly manipulating or disrupting these parameter vectors, specific knowledge can be edited or unlearned (hong2024intrinsic; hong-etal-2024-dissecting; meng2022locating), leading to changes in the model’s responses.
@@
elhage2022toy; olah2023superposition propose the concept of Knowledge Superposition. It refers to an inevitable phenomenon in neural network models, especially large language models, during training and data memorization: since the number of data features greatly exceeds the number of parameters in the model, each parameter does not have a simple one-to-one mapping with the data features or knowledge. Neurons are often involved with multiple data features simultaneously. In our work, we treat each vector in the last layer of MLP as a basic unit for storing knowledge and investigate the superposition of knowledge within these vectors.
@@ 3.1 Preliminary
In transformer-based language models, the MLP is a crucial component for storing the model’s factual knowledge, and its sub-layers can be viewed as key-value memories (geva2021transformer). To be specific, the first layer***In most decoder-only models, such as GPT-2 (radford2019language) and GPT-J (chen2021evaluating), the MLP component consists of two layers, whereas in LLaMA (touvron2023llama), it comprises three layers. However, we can still regard LLaMA’s first two layers collectively as the key matrices, with their output representing the coefficient scores. of MLP sublayers can be viewed as a matrix $W_{K}$ formed by key vectors $\{\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{n}\}$, used to capture a set of patterns in the input sequence, and ultimately outputting the coefficient scores. The second layer can be viewed as a matrix $W_{V}$ formed by value vectors $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}\}$, with each value vector containing the corresponding factual knowledge.
@@ 3.2 Knowledge Vectors Masking Procedure
Therefore, given a concept, when we aim to identify which specific value vectors in the model’s MLPs are most closely related to the knowledge contained in that concept—while avoiding the masking of vectors associated with the model’s general capabilitiesThe term ”general ability” refers to the model’s fundamental skills, such as processing text inputs correctly and generating coherent outputs, rather than encoding knowledge specific to a particular concept. (meng2022locating; geva2023dissecting), i.e., determining the appropriate subset $S^{\ell}$ at each layer of the model for this concept—we will run $t$ concept-related questions and $t^{*}$ irrelevant questions on the selected model. Then we will compute the corresponding coefficients $\mathbf{m}^{\ell}$ and $\mathbf{m}^{*\ell}$, which are the averages of the coefficients for the concept-related questions and irrelevant questions, respectively, at each layer of the model. For details on the generation of concept-related and irrelevant questions, as well as the selection of $t$ and $t^{*}$, please refer to §3.4. After obtaining $\mathbf{m}^{\ell}$ and $\mathbf{m}^{*\ell}$ at each layer, we perform the computation using the following formula:
$\mathbf{S^{\ell}}=\left\{\left|m^{\ell}_{j}-m^{*\ell}_{j}\right|\;\middle|\;1\leq j\leq n,\;m^{\ell}_{j}\in\mathbf{m}^{\ell},\;m^{*\ell}_{j}\in\mathbf{m}^{*\ell}\right\}$ (5)
@@ 3.3 The Definition of Parameter Specialization
[…]
@@ 3.4 Dataset Construction
Given that the frequency of knowledge in training datasets significantly influences a model’s ability to retain and comprehend it (allenzhu2023physics; meng2022locating; mallen-etal-2023-trust), we utilize Wikipedia page views as a proxy for knowledge frequency in the models’ pre-training datasetsTo better support this point, we include experiments in §A.4 of Appendix that validate the strong correlation between concept popularity and their frequency in the pretraining data.. To this end, we calculated the page views for each concept on Wikipedia between January 1, 2010, and December 31, 2019The earliest release date of the evaluated models, such as GPT-J, is 2019. Therefore, their pretraining datasets could not include knowledge or concepts that emerged after this period. To ensure a fair evaluation across all models, any knowledge introduced post-2019, including the COVID-19 pandemic, was excluded from the benchmark. . Based on these statistics, concepts are categorized by page view frequency into three equal tiers: low-frequency (bottom 33% of the distribution), medium-frequency (middle 33%), and high-frequency (top 33%). A more detailed distribution of the categories and the corresponding example data of SpecWiki dataset are provided in Table 4 and Table 5, respectively, in the Appendix. This approximation helps estimate the likelihood of a concept’s presence in the models’ pre-training datasets and allows us to explore how knowledge at different frequency levels is stored in models and provides a more comprehensive evaluation.
@@
  • Multi-Choice Questions. Drawing inspiration from the widely used Massive Multitask Language Understanding (MMLU) dataset (hendryckstest2021), which evaluates general knowledge across models, we similarly designed ten multiple-choice questions for each concept, ensuring the knowledge and answers could be directly found in the relevant Wikipedia articles. Specifically, we provided GPT-4o (openai2024gpt4ocard) with the appropriate Wikipedia article for each concept and instructed it to extract ten questions without overlap, along with the correct answers derived from the article’s text. Next, it was instructed to generate three additional incorrect answers, aside from the golden answer, ensuring that none of them overlapped with the correct answer to avoid confusion. The detailed prompt is available in §A.1. We also include sample multiple-choice questions and results of manual verification of the generated data in Appendix §A.2.
@@ 4.1 Experimental Setup
To provide a more comprehensive evaluation of how the degree of parameter specialization evolves across large language models, we assessed 20 open-source models from various families and sizes in the community. Specifically, we evaluated LLaMA series (touvron2023llamaopenefficientfoundation; touvron2023llama; grattafiori2024llama3herdmodels), Qwen series (bai2023qwentechnicalreport; yang2024qwen2technicalreport; qwen3blog), Gemma series (gemmateam2024gemmaopenmodelsbased; gemmateam2024gemma2improvingopen), OLMo series (groeneveld2024olmo; olmo20252olmo2furious), Yi series (ai2025yiopenfoundationmodels), Mistral series (jiang2023mistral7b), GPT-j-6b (wang2021gptj), Pythia-6.9b (biderman2023pythiasuiteanalyzinglarge), Falcon-7b (almazrouei2023falconseriesopenlanguage) and Mpt-7b (databricks2023mpt7b). Refer to Appendix §B.1 for the implementation details of these models.
@@
Regarding the selection of model layers for masking, since the initial layers of a model typically handle fundamental capabilities like basic text processing (meng2022locating; geva2023dissecting), masking these layers could severely impair the model’s basic text generation abilities. Therefore, for all models in our study, we preserve the first 5 layers without masking and only apply vector masking operations to all subsequent layers.
@@ 4.2 Main Results
The main results for the Multiple-Choice Questions setting can be seen in Figure 2. We observe a strong correlation between the degree of Parameter Specialization (measured by PSS) and model performance on SpecWiki across 20 models, with Pearson and Spearman coefficients of 0.92 and 0.93, respectively. Models achieving better performance on SpecWiki exhibit higher Parameter Specialization Scores. Furthermore, models with higher PSS are often those released more recently (warmer color) and exhibit stronger general abilities, as measured by their MMLU performance (larger circle). The corresponding results for the Open-ended Generation setting can be found in Figure 6 in §B.2, which exhibit similar patterns and trends.
@@
[…]
Qwen1.5-0.5B 0.61 $(\pm 0.2)$ 0.019 $(\pm 0.01)$
Qwen1.5-1.8B 0.61 $(\pm 0.3)$ 0.044 $(\pm 0.02)$
Qwen1.5-4B 0.73 $(\pm 0.2)$ 0.075 $(\pm 0.02)$
Qwen1.5-7B 0.75 $(\pm 0.2)$ 0.121 $(\pm 0.04)$
Qwen1.5-14B 0.82 $(\pm 0.2)$ 0.184 $(\pm 0.03)$
Gemma2-2B 0.72 $(\pm 0.3)$ 0.057 $(\pm 0.02)$
Gemma2-9B 0.86 $(\pm 0.1)$ 0.138 $(\pm 0.03)$
@@ 4.4 Evolution of Parameter Specialization During Pretraining
To better investigate the development of Parameter Specialization in the model from the perspective of model training dynamics, we analyzed 10 checkpoints from the OLMo-2-1124-7B (olmo20252olmo2furious) pretraining process by using our SpecWiki. The results are shown in Figure 4 below.
@@ 5 Validation of Parameter Specialization Benefits for Knowledge Tasks
[…]
LLaMA2-7B 0.60 $(\pm 0.2)$ 0.51 $(\pm 0.1)$ 0.67 $(\pm 0.1)$ 0.67 $(\pm 0.1)$ 11.23 $(\pm 2.1)$
LLaMA2-7BFT-FV 0.63 $(\pm 0.3)$ 0.54 $(\pm 0.2)$ 0.65 $(\pm 0.2)$ 0.62 $(\pm 0.1)$ 11.12 $(\pm 1.4)$
LLaMA2-7BFT-PV 0.67 $(\pm 0.3)$ 0.59 $(\pm 0.2)$ 0.72 $(\pm 0.1)$ 0.50 $(\pm 0.2)$ 7.89 $(\pm 2.1)$
LLaMA2-7BFT-CV 0.62 $(\pm 0.1)$ 0.51 $(\pm 0.1)$ 0.63 $(\pm 0.2)$ 0.62 $(\pm 0.1)$ 11.12 $(\pm 1.4)$
LLaMA2-7BFT-RV 0.58 $(\pm 0.2)$ 0.49 $(\pm 0.2)$ 0.65 $(\pm 0.2)$ 0.65 $(\pm 0.2)$ 11.07(±2.7)($\pm$ 2.7)
Qwen2-7B 0.72 $(\pm 0.3)$ 0.63 $(\pm 0.1)$ 0.124 $(\pm 0.03)$ 0.56 $(\pm 0.1)$ 9.78 $(\pm 1.9)$
Qwen2-7BFT-FV 0.73 $(\pm 0.2)$ 0.67 $(\pm 0.2)$ 0.110 $(\pm 0.02)$ 0.59 $(\pm 0.2)$ 8.53 $(\pm 1.1)$
Qwen2-7BFT-PV 0.77 $(\pm 0.1)$ 0.70 $(\pm 0.1)$ 0.133 $(\pm 0.03)$ 0.39 $(\pm 0.1)$ 6.92 $(\pm 1.3)$
Qwen2-7BFT-CV 0.73 $(\pm 0.2)$ 0.65 $(\pm 0.2)$ 0.114 $(\pm 0.03)$ 0.55 $(\pm 0.2)$ 8.78 $(\pm 1.6)$
Qwen2-7BFT-RV 0.71 $(\pm 0.2)$ 0.63 $(\pm 0.1)$ 0.122 $(\pm 0.02)$ 0.59 $(\pm 0.2)$ 9.65 $(\pm 1.4)$
@@
In this section, we conducted four sets of controlled training experiments, each involving continued fine-tuning on the Llama2-7B-base (touvron2023llamaopenefficientfoundation) and Qwen2-7B-base (yang2024qwen2technicalreport) models with additional knowledge data. These experiments aim to validate the causal relationship between improved parameter specialization and enhanced model performance on knowledge tasks.